Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

5 blue marbles, 3 red marbles and 4 purple marbles

Expert replies
by VJesus12 » Fri Apr 06, 2018 2:35 am
5 blue marbles, 3 red marbles and 4 purple marbles are placed in a bag. If 4 marbles are drawn without replacement, what is the probability that the result will not be 2 blue and 2 purple marbles?

A. 4/33
B. (5/36)^2
C. 1/2
D. (31/36)^2
E. 29/33

The OA is E.

Can someone show me what is the best approach to solve this PS question? Can it be solved in less than 5 minutes?
Join the discussion
Source: — Problem Solving |

by GMATGuruNY » Fri Apr 06, 2018 5:55 am
VJesus12 wrote:5 blue marbles, 3 red marbles and 4 purple marbles are placed in a bag. If 4 marbles are drawn without replacement, what is the probability that the result will not be 2 blue and 2 purple marbles?

A. 4/33
B. (5/36)^2
C. 1/2
D. (31/36)^2
E. 29/33
P(not 2 blue and 2 purple) = 1 - P(exactly 2 blue and exactly 2 purple).

P(exactly n times) = P(one way) * total possible ways.

P(one way):
One way to get exactly 2 blue and exactly 2 purple:
BBPP.
P(B on the 1st pick) = 5/12. (Of the 12 marbles, 5 are blue.)
P(B on the 2nd pick) = 4/11. (Of the 11 remaining marbles, 4 are blue.)
P(P on the 3rd pick) = 4/10. (Of the 10 remaining marbles, 4 are purple.)
P(P on the last pick) = 3/9. (Of the 9 remaining marbles, 3 are purple.)
Since we want all of these events to happen, we MULTIPLY:
5/12 * 4/11 * 4/10 * 3/9 = 2/99.

Total possible ways:
BBPP is only ONE WAY to get exactly 2 blue and exactly 2 purple.
Now we must account for ALL OF THE WAYS to get exactly 2 blue and exactly 2 purple.
Any arrangement of the letters BBPP represents one way to get exactly 2 blue and exactly 2 purple.
Thus, to account for ALL OF THE WAYS to get exactly 2 blue and exactly 2 purple, the result above must be multiplied by the number of ways to arrange the letters BBPP.
Number of ways to arrange 4 elements = 4!.
But when an arrangement includes IDENTICAL elements, we must divide by the number of ways each set of identical elements can be ARRANGED.
The reason:
When the identical elements swap positions, the arrangement doesn't change.
Here, we must divide by 2! to account for the 2 identical B's and by another 2! to account for the 2 identical P's:
4!/(2!2!) = 6.

Multiplying the results above, we get:
P(exactly 2 blue and exactly 2 purple) = 2/99 * 6 = 4/33.
Thus:
P(not 2 blue and 2 purple) = 1 - 4/33 = 29/33.

The correct answer is E.

More practice:
https://www.beatthegmat.com/select-exact ... 88786.html
https://www.beatthegmat.com/probability- ... 14250.html
https://www.beatthegmat.com/a-single-par ... 28342.html
https://www.beatthegmat.com/at-a-blind-t ... 20058.html
https://www.beatthegmat.com/rain-check-t79099.html
https://www.beatthegmat.com/probability-t227448.html
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by Scott@TargetTestPrep » Mon Apr 16, 2018 4:16 pm
VJesus12 wrote:5 blue marbles, 3 red marbles and 4 purple marbles are placed in a bag. If 4 marbles are drawn without replacement, what is the probability that the result will not be 2 blue and 2 purple marbles?

A. 4/33
B. (5/36)^2
C. 1/2
D. (31/36)^2
E. 29/33

We can use the formula:

P(will not be 2 blue and 2 purple marbles) = 1 - P(2 blue and 2 purple marbles)

We can select two blue marbles in 5C2 = (5 x 4)/2! = 10 ways.

We can select 2 purple marbles in 4C2 = (4 x 3)/2! = 6 ways.

Thus, the total number of pairings of 2 blue and 2 purple marbles is 10 x 6 = 60 ways.

We can select 4 marbles from 12 in 12C4 = (12 x 11 x 10 x 9)/(4 x 3 x 2) = 11 x 5 x 9 = 495 ways.

Thus, P(2 blue and 2 purple marbles) = 60/495 = 20/165 = 4/33.

So, P(will not be 2 blue and 2 purple marbles) = 1 - 4/33 = 29/33.

Answer:E

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion